OpenStudy (anonymous):

Two I-beams are to be joined end to end. One has been measured to be 2.680 m long; the other, 4.9 m. When they are bolted together, what value can you accurately give for the total length? I. Tenths (7.6 m) II. None of these III. Whole number (8 m) IV. Hundredths (7.58 m) V. Two significant figures (7.6 m) VI. Four significant figures (7.580 m) VII. Three significant figures (7.58 m)

OpenStudy (anonymous):

I think the answer is "II. None of these"

OpenStudy (anonymous):

But I'm not sure.

OpenStudy (shane_b):

For addition and subtraction the result should have as many decimal places as the measured number with the smallest number of decimal places. So what does that tell you?

OpenStudy (anonymous):

When add these two together, I got 7.58, but I know what answer to choose.

OpenStudy (anonymous):

When I add these two together, I got 7.58, but I know what answer to choose.*

OpenStudy (shane_b):

Ok so you have 7.58 as the answer. Now which of the two measured values has the smallest number of decimal places: 4.9 or 2.680?

OpenStudy (anonymous):

4.9

OpenStudy (shane_b):

So your answer should only have 1 decimal place (tenths). So the answer can be rounded to the tenths place.. 7.6m

OpenStudy (theeric):

Ah, didn't see other replies! Shane_B will make sure you have this. Here's a link for more: http://chemed.chem.purdue.edu/genchem/topicreview/bp/ch1/sigfigs.html

OpenStudy (anonymous):

So, the final answer will be "Tenths (7.6 m), right?

OpenStudy (shane_b):

Yes

OpenStudy (anonymous):

You guys are so good! Thank you so much! However, I would like, can you still help me with another question?

OpenStudy (shane_b):

Yes, post it separately though so others can help if needed

OpenStudy (anonymous):

Okay, I will right! Thank you!

OpenStudy (shane_b):

np :)

OpenStudy (anonymous):

Posted it! :)

OpenStudy (anonymous):

Consider the equations I. v = v0 + ax II. y = (2m) cos(kx), where k = 2m^-1 III. v^2 = ax Which ones are dimensionally correct?